OpenStudy (anonymous):

The graph of a line has a slope of 2 and passes through the point (–6, –5). What is the equation of the line in point-slope form?

OpenStudy (anonymous):

\[y-y _{1}=m \left( x-x _{1} \right)\] then simplify it.

OpenStudy (anonymous):

so y + 6 = 2(x + 5)?

OpenStudy (anonymous):

\[x _{1}=-6,y _{1}=-5\]

OpenStudy (anonymous):

the other the way around?

OpenStudy (anonymous):

y + 5 = 2(x + 6)